OpenStudy (raiders88):

multiply the binomials (3x-2)(3x+2)

OpenStudy (gabbyalicorn):

(3x−2)(3x+2) =(3x+−2)(3x+2) =(3x)(3x)+(3x)(2)+(−2)(3x)+(−2)(2) =9x2+6x−6x−4 =9x2−4

OpenStudy (anonymous):

answer choices?

OpenStudy (raiders88):

gabbay i have a few more

OpenStudy (gabbyalicorn):

Okay.

OpenStudy (gabbyalicorn):

Just please make a new question. Thanks.

OpenStudy (raiders88):

ok